The equation of the ellipse is given as "(x\/a)^2+(y\/b)^2 =1"

find the equation of the tangent at (x0, y0)


1
Expert's answer
2021-07-26T16:31:00-0400

"\frac{d}{dx}((\frac{x}{a})^2+(\frac{y}{b})^2)=\frac{d}{dx}(1)."

"2\frac{x}{a^2}+2\frac{y}{b^2}y'=0."

"y'=-\frac{b^2}{a^2}\frac{x}{y}."

"y'(x_0,y_0)=-\frac{b^2}{a^2}\frac{x_0}{y_0}."

Tangent line: "y-y_0=-\frac{b^2}{a^2}\frac{x_0}{y_0}(x-x_0)."

Or "y=-\frac{b^2}{a^2}\frac{x_0}{y_0}x+\frac{b^2}{a^2}\frac{x_0}{y_0}x_0+y_0."